6) Using the copied portion of the Results section, use these values: (hint: MSE is a common notation for MSerror; also you might find it helps to work backward from the F value that is given)

SS

df

MS

F

p

Request Type

error

total

For low Rational scorers,

donations differed significantly by request type, F(2, 55) = 3.35, p < .05,

η2 = .109, MSE. 2.98. In planned comparisons based on Small et al.’s

findings (2007, Study 3), donations were higher when a personalized victim

appeared alone (M = $3.18, n =17) than when statistics were added to the

victim request (M = $2.00, n = 22, p = .04) or when statistics were presented

alone (M = $1.79, n = 19, p = .02). The latter two means did not differ.
Answer #1

n=total no. of observations=17+22+19=58

Source SS df MS F p
Request Type 9.983*2=19.966 3-1=2 3.35*2.98=9.983 3.35 P(F>3.35|F~F2,55)=0.0424
Error 2.98*55=163.9 58-3=55 2.98
Total 20.1+163.9=184 58-1=57